OpenStudy (anonymous):

Write the equation of the line that contains (1,4) and (3,10) in 3 forms. Point slope, slope intercept, and standard form.

OpenStudy (lgbasallote):

huh?

OpenStudy (lgbasallote):

what does the last sentence mean?

OpenStudy (anonymous):

first you need the slope do you know how to find it? use \(m=\frac{y_2-y_1}{x_2-x_1}\) with \(x_1=3, x_2=1, y_2=10,y_1=4\) you get \[m=\frac{10-4}{3-1}=\frac{6}{2}=3\]

OpenStudy (anonymous):

then for the "point slope" form write \(y-y_1=m(x-x_1)\) \(m=3\) and you have a choice of points to use either (1,4) or (3,10) i choose (3,10) so you get \(y-10=3(x-3)\)

OpenStudy (anonymous):

then "slope intercept " form means take the above equation and solve for \(y\)

OpenStudy (anonymous):

"standard form" means write it as \(ax+by=c\)